# Understanding the Local Geometry of Generative Model Manifolds

### v1

@@ 4 Guiding Generation With Geometry
Reward Modeling. In the previous chapters, we established that geometric descriptors are correlated with out-of-distribution detection, memorization, and aesthetics. Our primary focus has been on analyzing the geometry of the learned data manifold. In this section, we demonstrate that these geometric descriptors can be utilized to intervene in the latent representations of generative models, resulting in meaningful effects on the generated examples. Leveraging recent instance-level guidance methods, such as those described in the Universal Guidance paper [2], we can effectively influence the latents to produce significant alterations in the generated outputs. To guide the generation process, we employ local scaling as our geometric descriptor. We use classifier guidance to perform gradient ascent with step size $\rho$ on gradients derived from local scaling. This method allows us to steer the generative process by maximizing the local scaling descriptor, thus altering the generated examples in a controlled manner.
@@
Figure 10: Reward guidance on stable diffusion (maximizing the reward). Local scale reward guidance increases from left to right in each picture, with the first image showing no reward guidance. We observe maximizing the reward leads to sharper details, improved sharpness and contrast, and higher diversity in the images. Prompts used for generation from top to bottom: “a Rhodesian ridgeback with a city in the background”, “a Beagle in the snow”, “a Australian terrier on a cobblestone street”
Figure 11: Reward guidance on stable diffusion (minimizing the reward). Local scale reward guidance decreases from left to right in each picture, with the first image showing no reward guidance. We observe that decreasing the local scale leads to minimized uncertainty, resulting in a noticeable blurring effect and loss of details in the image. Prompts used for generation from top to bottom:“a Beagle with a city in the background”, “a Samoyed with a mountain in the background”, “a Australian terrier on a cobblestone street”.
To enable universal guidance, we need to compute the gradient of the reward function. In our context, this entails calculating the Hessian matrix, which is computationally expensive. Instead, we train a reward model to approximate local scaling, given noisy latents at time step $t$ of the diffusion chain. We transform the regression task of estimating a continuous local scaling value into a classification task by first determining the maximum and minimum range of local scaling values for the ImageNet dataset. We then discretize this range into five uniform bins. Our training data collection process is as follows: For each image in the ImageNet dataset, we sample 10 time steps, apply noise to the encoded latents according to the sampled time step, and compute the corresponding local scaling. Subsequently, we train a classification model on pairs of latents and their associated local scaling bins.
@@
Our experiments reveal that maximizing local scaling in the manifold of a stable diffusion model directly correlates with adding texture to the generated images. Moreover, this approach enhances diversity based on single images. By optimizing the local scaling descriptor, the generative model is guided towards producing more varied and textured outputs. This approach is notable because traditional methods for diversity guidance generally function at the distribution level. Our method, however, focuses on maximizing the inherent diversity as preserved by the model within its learned manifold, effectively steering the generated images towards the extremities of the distribution. This instance-level intervention allows for a more detailed and precise enhancement of diversity, presenting a novel approach to guiding generative models.
@@
As seen from Fig. 10 maximizing the reward results in added details in form of sharpening the image, adding texture and contrast. We also observe that if we move towards minimizing the reward, the images tend to loose fine-grained details as seen in Fig. 20.
@@ 5 Conclusion & Future Directions
In this paper, we proposed a novel self-assessment approach to evaluate generative models using geometry-based descriptors – local scaling ($\psi$), local rank ($\nu$) and local complexity ($\delta$) - effectively while utilizing only the model’s architecture and weights. Our approach characterizes uncertainty, dimensionality, and smoothness of the learned manifold without requiring original training data or human evaluators. Our experiments demonstrated how these descriptors relate to generation quality, aesthetics, diversity, and biases. We showed that the geometry of the data manifold impacts out-of-distribution detection, model comparison, and reward modeling, enabling better control of output distribution. While using the geometry of manifolds offers a novel approach to self-assess generative models, we acknowledge two main limitations that warrant further investigation. First, the geometry of the learned manifold is inherently influenced by the training dynamics of the model. A deeper understanding of this relationship is needed to fully leverage geometric analysis for model assessment and improvement. Second, the computational complexity of our method, particularly the calculation of the Jacobian matrix, poses a practical challenge, especially for large-scale models. Future work should explore more efficient algorithms or approximations to address this limitation.
Appendix A Appendix / Supplemental material
Appendix B Broader Impact Statement
Appendix C Extra Figures
Local scaling (log), $\psi$
Figure 13: Sub-population differences of local descriptors in generated data. Separation between the $\psi$ distributions for $1000$ images generated for each prompt in a photo of a {rich, poor} swiss person (left) and a photo of a {bartender, crane operator} (right). We see that the uncertainty as expressed by $\psi$ can vary between sub-populations indicating an underlying bias in generation diversity.
Low local rank, $\nu\downarrow$
High local rank, $\nu\uparrow$
Figure 14: Images with the lowest (left) and highest (right) local rank $\nu$ from a set of $20000$ randomly sampled ImageNet dataset samples. Low rank images contain simpler textures for every class compared to the high rank samples. This is because for images with higher local rank, the learned manifold is higher dimensional therefore allowing higher independent degrees of variations locally for the generated images.
Low local scaling, $\psi\downarrow$
High local scaling, $\psi\uparrow$
Figure 15: Imagenet images with high and low local scaling for the stable diffusion decoder. Each coordinate in both left and right image grids, correspond to the same imagenet class.
Figure 16: Bivariate $(\epsilon,\psi)$ and $(\epsilon,\nu)$ distributions for training samples in green (top), and $(\delta,\psi)$, $(\nu,\psi)$ distributions for training in green and generated samples in red (bottom). Local complexity $\delta$ exhibits linear correlation with $\epsilon$ and $\psi$. Lowest $\epsilon$ samples also have lowest $\delta$, indicating that good reconstruction requires linear interpolation of the data manifold. Local rank $\nu$ uncorrelated with $\epsilon$. Both $\delta$ and $\nu$ exhibit some linear relationship with $\psi$. Note that the modes of real and generated samples for the $(\psi,\delta)$ and $(\psi,\nu)$ distributions are not completely aligned.
Figure 17: Local scaling dynamics while fine-tuning a latent diffusion model on Chest-Xray images. The local scaling is computed for generated chest-xray images, where we can see a reduction of local scaling, i.e., uncertainty.
@@ Appendix C Extra Figures
Figure 18: Geometric descriptors of the data manifold. Level sets of data manifold descriptors for a Beta-VAE trained unconditionally on MNIST. From left to right, we present training samples (top row) and generated samples (bottom row) for linearly increasing level sets of local scaling ($\psi$) from $[-80,-42]$, local complexity ($\delta$) from $[0,120]$ and local rank ($\nu$) from $[1.5,5.5]$. Not all level sets had an equal number of samples from training/generated distributions. We see that for higher $\psi$, we have more outlier samples whereas for lower $\psi$ we have modal samples. For increasing $\delta$ we see that the quality of generated samples decreases and the diversity of samples is reduced as well. For higher $\nu$ digits become more regularly shaped.

### v2

@@ 5 Guiding Generation With Geometry
In the previous sections, we have presented qualitative and quantitative evidence, establishing the connection between geometric descriptors and downstream generation. Among the three descriptors we find that local complexity has the highest sensitivity to aesthetic changes in images due to corruptions (fig. 6), and correlates with visual complexity (fig. 17). We also observe in fig. 5 higher local scaling level sets for samples generated using a classifier free guidance scale of $7.5$, have higher diversity while maintaining higher predicted human preference scores. Based on these results, we wish to explore whether local scaling can be used to guide generation.
Local Scaling
Vendi Score
Hum. Pref. Score
Pre-guidance local scaling level set $\rightarrow$
Figure 8: Guidance via reward models trained with increasing number of training samples. We observe that even with when only 50K samples are used in training, reward models can increase generation diversity, local scaling and human preference scores. Here we present the change of vendi score, average local scaling and average aesthetic score for pre-guidance local scaling level sets increasing from left to right on the x-axis.
Recently proposed instance-level universal guidance method (Bansal et al., 2023), can effectively influence the latents in the reverse process of a latent diffusion models to produce desired changes. Directly using local scaling to guide generation using such methods, require calculating the input-output Hessian since local scaling is a first-order measure that requires computing the input-output jacobian. To avoid computing the Hessian we train a reward model as a proxy and use the reward model gradients directly. Instead of training on continuous local scaling values in a regression task, we transform it into a local scaling level set classification task. We discretize the range of local scaling values into $5$ bins and use the bin indices as training labels.
a Australian terrier with
a city in the background
$\rho=0$     $\rho=1$     $\rho=2$
a Australian terrier
on a cobblestone street
$\rho=0$     $\rho=1$     $\rho=2$
a Samoyed with a
mount. in the bkg.
$\rho=0$     $\rho=-1$     $\rho=-3$
a Beagle with
a city in the background
$\rho=0$     $\rho=-1$     $\rho=-3$
@@
Figure 9: Reward guidance on stable diffusion. Local scaling reward guidance increases from left to right in each picture (left-panel) and decreases left to right in each picture (right-panel), with the first image showing no reward guidance. We observe maximizing the reward leads to sharper details, improved sharpness and contrast, and higher diversity in the images. Decreasing the local scaling leads to minimized uncertainty, resulting in a noticeable blurring effect and loss of details especially in the background of the image.
Data preparation. We obtain training data for the reward model by i) sampling $N$ images from Imagenet and encoding them to the Stable Diffusion latent space ii) adding noise using the forward diffusion process up to randomly chosen noise levels iii) for each latent computing the local scaling descriptor.
To evaluate the performance of the reward model and dependency of the reward model on the number of training samples, we train multiple models for $N=$ $50K,200K,400K,800K$ For evaluation, we generate $2560$ samples using the dreambooth live subject prompt templates Ruiz et al. (2023), with Imagewoof Howard (2019) dogs as subjects. While Imagewoof dog classes are present in Imagenet therefore possibly in the training data, the dreambooth prompt templates contain a variety of settings that are not generally present in Imagenet, e.g., ‘a <subject> on top of pink fabric’.
Evaluation Setup. We first sample Stable diffusion without any reward guidance and with classifier-free guidance of 7.5 to obtain baseline samples. We partition the range of local scaling values obtained for the baseline samples into $n=10$ bins LABEL:{fig:reward_model_vendi_aesthetics}, where each bin contains images from a local scaling level set. Following that we use the same seed and prompts as the baseline samples to generate images using reward guidance to increase local scaling. For each bin or pre-guidance local scaling level set, we compare between the baseline samples and corresponding reward guided generations in the following three axes: i) change of local scaling ii) change of vendi (diversity) score iii) the change of human preference score (RAHF Liang et al. (2024) aesthetic score).
Results. In fig. 8, we present the mean local scaling per bin with 95% confidence interval. Here the blue line represents the mean pre-guidance local scaling values, increasing from left to right. In LABEL:{fig:reward_model_vendi_aesthetics}, we present vendi scores and average predicted human preference scores. For any bin, we present results for the reward guidance scale that maximizes the local scaling. We see that even for a model trained with $50K$ samples, we can have a considerable increase in the local scaling, diversity and aesthetic score for most of the bins. Changes in local scaling, vendi and aesthetic scores are higher for the lower pre-guidance local scaling level set bins compared to the higher pre-guidance local scaling level set bins.
@@
Our experiments reveal that maximizing local scaling in the manifold of a stable diffusion model directly correlates with adding texture to the generated images. Moreover, this approach reduces the likelihood on the manifold for single images. By optimizing the local scaling descriptor, the generative model is guided towards producing more varied and textured outputs.
This approach is notable because traditional methods for diversity guidance generally function at the distribution level. Our method, however, focuses on maximizing the inherent diversity as preserved by the model within its learned manifold, effectively steering the generated images towards the extremities of the distribution. This instance-level intervention allows for a more detailed and precise enhancement of diversity, presenting a novel approach to guiding generative models.
@@
As seen from Fig. 9 (left-panel) maximizing the reward results in added details in form of sharpening the image, adding texture and contrast. We also observe that if we move towards minimizing the reward, the images tend to loose fine-grained details as seen in Fig. 9 (right-panel). Please refer to the supplementary material for more visual results.
@@ 6 Conclusion & Future Directions
In this paper, we present empirical evidence that the local geometric descriptors – local scaling ($\psi$), local rank ($\nu$) and local complexity ($\delta$) - can effectively characterize the local geometry and distinguish between downstream qualitative aspects of generated samples such as generation quality, aesthetics, diversity, and memorization. Such descriptors only utilize the model’s architecture and weights to characterize the behavior of generative models. We acknowledge two main limitations that warrant further investigation. First, the geometry of the learned manifold is inherently influenced by the training dynamics of the model. A deeper understanding of this relationship is needed to fully leverage geometric analysis for models. Second, the computational complexity of our method, particularly the calculation of the Jacobian matrix, may pose a practical challenge, especially for large-scale models. Future work should explore more efficient algorithms or approximations to address this limitation.
Appendix A Computation times for local descriptors
The computation times for the local scaling and local rank computation (since both require one randomized SVD computation for one latent vector) ends up being 3929s for 1000 samples. For local complexity we require 113s for 1000 samples. All the estimates are for a JAX implementation of Stable Diffusion on TPUv3.
Note that to train a reward model, we require the descriptors to be computed only once for each pre-trained model. If we compute the local scaling for 100k samples we require 173.1 TPU v3 hours which is equivalent to 54.58 V100 hours (according to Appendix A.3 Dhariwal & Nichol (2021)). Compared to 79,000 A100 hours required for Stable Diffusion training [^2] https://www.mosaicml.com/blog/training-stable-diffusion-from-scratch-costs-160k,  24000 hours with enterprise level optimization [^3] https://www.databricks.com/blog/stable-diffusion-2, the computation required for the descriptors and reward model training is significantly small. The computation time for the local descriptors can be further reduced by using a smaller $k$ for our projection matrix $W$, or by using non-jacobian based methods, e.g., estimating the local scaling by measuring the change of volume for a unit norm $\ell_{1}$-ball in the input space. We leave exploration of these directions for future work.
Appendix B Related Works
Local geometry pre-diffusion. Early applications of the local geometry of generative models involved improving the generation performance and/or utility of generative models via geometry inspired methods. For example, in Rifai et al. (2011) the authors proposed regularizing the contraction of the local geometry to learn better representations in autoencoders trained on MNIST and CIFAR10. The regularization penalty is employed via the norm of the input-output jacobian in Rifai et al. (2011), is an upper bound for local scaling presented in our paper. In Arvanitidis et al. (2017) the authors provided visualizations on the curvature of pre-trained VAE latent spaces and proposed using an auxiliary variance estimator neural network to regularize the latent space geometry during generation. In Kuhnel et al. (2018) the authors perform latent space statistical inference problems, e.g., maximum likelihood inference, by training a separate neural network to approximate the Riemannian metric. In Humayun et al. (2022a) the authors proposed a novel latent space sampling distribution based on the latent space geometry that allows uniformly sampling the learned data manifold of continuous-piecewise affine generators. The authors showed downstream benefits with fairness and diversity for such latent space samplers. While most of these methods discuss pre-diffusion architectures, their results are early demonstrations of how the local geometry can affect downstream generation. also employ auxiliary Neural Networks to model an intrinsic property of a pre-trained generator, similar to how we propose using a reward model for Stable Diffusion.
Local intrinsic dimensionality of diffusion models. The local geometry of diffusion models and possible applications have garnered significant interest in recent years. In stanczuk2022your the authors propose a method to compute the intrinsic dimensionality of diffusion models using the assumption that the score field is perpendicular to the data manifold. For any vector $x$ on the data manifold, the method requires computing the dimensionality of the score field around $x$ and subtracting it from the ambient dimension. To do that, the authors perform one step of the forward diffusion process $k$ times for $x$, denoise the $k$ noisy samples using the diffusion model and compute the rank of the data matrix containing denoised samples to obtain the intrinsic dimensionality. Compared to this method, we compute the dimensionality directly via a random estimation of the input-output jacobian SVD. We do not require any assumption on the score function vector field being perpendicular to the data manifold, which may not hold for a diffusion model that is not optimally trained or highly complex training datasets like LAION.
In Kamkari et al. (2024) the authors compute rank using the method proposed in stanczuk2022your and show that local intrinsic dimensionality can be used for out-of-distribution (OOD) detection. This is analogous to our analysis in Sec 3 on the local geometry on or off the manifold. We can see that the intuition authors provided in Kamkari et al. (2024) for diffusion models trained on smaller models and datasets e.g., FMNIST, MNIST, translate to larger scale models like Stable Diffusion trained on LAION as we have presented fig. 4, fig. 17 and Sec 4. Especially in fig. 6, we show that creating OOD samples with corruptions on Imagenet data (in-distribution), we can have an increase or decrease in negative-log likelihood (estimated via local scaling), with decrease for blurring corruptions and increase in noising corruptions.
Concurrent work kamkari2024efficient has also shown the relationship between the intrinsic dimensionality (local rank) of Stable Diffusion scale models and the texture/visual complexity of generated images. We believe our analysis is much more holistic with three different geometric properties being measured compared to only local dimensinality. We i) show quantitatively how diversity measured via vendi score is higher for higher local scaling and rank values (fig. 5). We have explored how rank and scaling evolves continuously across the latent space in fig. 4. We have presented how the geometry distribution varies as we continually perturb images via noise or blurring operations fig. 6 And finally in Sec 5 we have presented a method to guide generation using the local geometry to obtain downstream generation benefits.
Misc. Apart from the aforementioned works, Kadkhodaie et al. (2023) show that the emergence of generalization in diffusion models – when two networks separately trained on the same data learn the same mapping – can be attributed to the eigenspectrum and eigenvectors of the input-output jacobian. While we do not study the training dynamics of the local geometric descriptors in our paper, Kadkhodaie et al. (2023) suggests that the local geometry can be an important indicator of memorization and generalization emergence in diffusion models. In Manor & Michaeli (2023) the authors use the posterior principal components of a denoiser for uncertainty quantification. This work suggests that components with larger eigenvalues result in larger uncertainty which is directly related to the local scaling descriptors as it measures the product of non-zero singular values. While in Manor & Michaeli (2023) the authors propose using it for only a single image denoiser, we show that it generalizes for any diffusion model including Stable Diffusion scale text-to-image diffusion models.
Appendix C Correlations between the three descriptors
Local scaling characterizes the change of volume by the affine slope $\mathbf{A_{\omega}}$ going from the latent space to the data manifold. Local rank characterizes the number of dimensions retained on the manifold after the network locally scales the latent space. Both local rank and scaling quantify first order properties of the CPWL operator. Local complexity approximates the ‘number of unique affine maps’ within a given neighborhood Humayun et al. (2024) by computing the number of CPWL knots intersecting an $\ell_{1}$ ball in the input/latent space. Therefore local complexity is a measure of ‘un-smoothness’ and quantifies local second-order properties of a CPWL operator.
Correlations between local scaling $\psi$ and local rank $\nu$. By definition, local scaling and local rank are correlated, since both characterize the change of volume by the network input-output map at any input space linear region – also evident in eq. 2 and eq. 4. Local scaling is also upper bounded by local rank, $\psi_{\omega}\leq\sigma_{0}^{\nu_{\omega}}$ where $\sigma_{0}$ is the largest singular value of $\bm{A}_{\omega}$. The correlation is evident for our low dimensional DDPM setting presented in fig. 3, local rank and local scaling are highly correlated in their spatial distribution. There are indications suggesting that the correlations persists throughout training as can be seen in fig. 3 rightmost column top and bottom. However in fig. 4, we can see that in the high-dimensional Stable Diffusion latent space, local scaling and rank are correlated but local rank has sharper changes spatially compared to local scaling.
Correlations between local complexity $\delta$ and rank $\nu$. There also exist correlations between local complexity and local rank due to the continuity of CPWL maps – between two neighboring linear regions $\omega_{1}$ and $\omega_{2}$, the corresponding slope matrices $\mathbf{A_{\omega_{1}}}$ and $\mathbf{A_{\omega_{2}}}$ differ by at most one row. Therefore between two neighboring regions $\omega_{1}$ and $\omega_{2}$, $|\nu_{\omega_{1}}-\nu_{\omega_{2}}|<=1$. Informally, the local rank in a neighborhood $V$ is lower bounded by the number of non-linearities in neighborhood $V$. This is evident in the empirical results presented in fig. 3 and fig. 4. In both figures, for input space neighborhoods with higher local complexity, we see a decrease in local rank. However, we do not observe sharp changes in local complexity as we observe in local rank in fig. 4. In fig. 3 we see that local rank is more discriminative of the data manifold compared to local complexity. Their training and denoising dynamics differ significantly as seen in fig. 3 rightmost column.
Qualitative and quantitative results on correlations. We train a beta-VAE unconditionally on MNIST and present in Fig. 10 samples from increasing local descriptor level sets from left to right along the columns. In Fig. 11, we present joint distributions of local scaling, complexity, rank and mean squared reconstruction error for training and test samples. We see that while local scaling, complexity and rank have some linear correlation, the classwise distribution in fig. 11 is very different between the three. We also present in fig. 12 the vendi score for increasing local scaling level sets and evidence that the population means for the descriptors don’t follow the same pattern between sub-populations.
Figure 10: Joint distributions for local scaling and MSE, local rank and MSE, local complexity and MSE, local scaling and local complexity, and local scaling and local rank. We observe that local complexity is linearly correlated wth MSE, with higher complexity images incurring higher error. Local scaling, rank and complexity have correlations between them as well.
@@ Appendix C Correlations between the three descriptors
Figure 11: Level sets of data manifold descriptors for a Beta-VAE trained unconditionally on MNIST. From left to right, we present training samples (top row) and generated samples (bottom row) for linearly increasing level sets of local scaling ($\psi$) from $[-80,-42]$, local complexity ($\delta$) from $[0,120]$ and local rank ($\nu$) from $[1.5,5.5]$. Not all level sets had an equal number of samples from training/generated distributions. We see that for higher $\psi$, we have more outlier samples whereas for lower $\psi$ we have modal samples. For increasing $\delta$ we see that the quality of generated samples decreases and the diversity of samples is reduced as well. For higher $\nu$ digits become more regularly shaped.
$\psi$
$\delta$
$\nu$
$\psi$
$\delta$
$\nu$
Figure 12: (Left panel) Vendi score (Friedman & Dieng, 2023) calculated for samples from different local descriptor level sets of a Beta-VAE. We take upto $150$ samples from each level set and compute vendi score seperately for the MNIST train dataset, test dataset and generated samples. (Right panel) Sub-population differences of local descriptors in training data. We see that the order of sub-population means for the three classes, are not the same for all three descriptors.
Appendix D Additional Experiments
D.1 Local scaling for transformer based diffusion model
Since we are based on the CPWL formulation of NNs, our framework would generalize to models of any scale and any architecture with CPWL non-linearities. Empirically we have shown it to generalize for non-CPWL architectures like Stable Diffusion v1.4 and DDPM that employs non CPWL non-linearities such as attention, GeLU and much more. We have performed additional experiments with a DiT-XL Peebles & Xie (2023) trained on Imagenet-256. For the DiT we compute the descriptors for the transformer network, conditioned on noise level $t=0$ , i.e., zero noise level. We generate 5120 images conditioned on Imagewoof Howard (2019) classes and present in fig. 13, increasing local scaling level sets from left to right. We see that similar to fig. 17 from the, DiT exhibits a qualitative correlation between visual complexity and local scaling. For additional analysis we repeat the Stable Diffusion experiments on the relation between diversity and local scaling for DiT. We see that similar to Stable Diffusion, for increasing local scaling level sets, the diversity of images increase and then drop for the highest local scaling level sets.
Figure 13: Left: Vendi score and membership counts for increasing local scaling level sets, computed for a DiT transformer. We see that similar to Stable Diffusion, local scaling increases from lower to higher local scaling level sets, then drops for very high local scaling level sets. Right: Generated samples from each level set in the left panel. Sample sets from higher local scaling level sets, tend to be more diverse.
$t=0.66T$
$t=0.44T$
$t=0.22T$
$t=0$
$\mathbb{E}_{\mathcal{M}}[\Phi]-\mathbb{E}_{\bar{\mathcal{M}}}[\Phi]$
$\psi^{t}$
$\delta^{t}$
$\nu^{t}$
Figure 14: Local geometric descriptors computed over the input domain of a pre-trained toy diffusion model trained to produce samples from a dinosaur manifold $\mathcal{M}\in\mathbb{R}^{2}$. Descriptors are computed by conditioning the diffusion model on noise level $t$. We consider the set of input vectors within $0.05$ units of the training data as on manifold $\mathcal{M}$ and rest as off the manifold $\bar{\mathcal{M}}$. We present the difference between the expected descriptor values on and off the manifold, $\mathbb{E}_{\mathcal{M}}[\Phi]-\mathbb{E}_{\bar{\mathcal{M}}}[\Phi],\forall\Phi\in\{\psi^{t},\delta^{t},\nu^{t}\}$ at different training iterations (right). We also present the descriptor computed over $[-6,6]^{2}$ for different noise levels $t$ after $125000$ training iterations (rest). We observe that $\psi^{t}$ is lower, $\delta^{t}$ is higher and $\nu^{t}$ is lower on the manifold than off the target manifold for lower noise levels, especially after the model is considerably trained. This indicates that for well trained diffusion model, i.e., learned manifold $\hat{\mathcal{M}}\approx\mathcal{M}$, local descriptors can distinguish between on and off manifold vectors in the input space.
D.2 VAE training dynamics for MNIST
$\delta$
$\psi$
optimization steps
[…]
Noise Level $t=[0.5,1]$
Figure 16: Local scaling distribution difference between in-domain (blue) vs out-of-domain (red) datasets when conditioned on different noise levels for an SD Unet trained on the CelebAHQ dataset. Here $\mathcal{T}(a,b)$ denotes a t-test between local scaling distributions for dataset $a$ and dataset $b$.
Observations. By increasing the noise we control the puffiness of the target manifold. We observe that as the noise standard deviation is increased there is 1) increase in $\delta$ indicating the manifold becomes less smooth 2) decrease in local scaling indicating that the uncertainty decreases. We can also observe an initial dip in both local complexity and local scaling. This is similar to what was observed for discriminative models in (Humayun et al., 2024) where a double descent behavior was reported in the local complexity training dynamics of classification models. Based on these results, contrary to the observation in (Humayun et al., 2024), generative models do not have a double descent in local complexity however we do observe a double ascent in local scaling. Our observations suggest that the training dynamics need to be taken into account, when comparing the local manifold geometry between two separately trained models.
Appendix E Entropy difference between two nearby region
Suppose we have an injective $\mathcal{G}:Z\rightarrow X$ mapping learned by a CPWL generator $\mathcal{G}$. Any linear region $\omega$ in the latent space CPWL partition $\Omega$ is mapped to a unique region on the output manifold. We define $S$ as:
$S=$ $\mathcal{G}(\mathbf{z})\forall\mathbf{z}\in\omega$ $=$ $\mathbf{A_{\omega}}$$\mathbf{z}$ +$\mathbf{b_{\omega}}$ $\forall\mathbf{z}\in\omega$
The change of volume from $\omega\rightarrow S$ is $\sqrt{det(\mathbf{A_{\omega}}^{T}\mathbf{A_{\omega}})}$. Therefore for any latent $z$ and output $x=\mathcal{G}(\mathbf{z})$:
$p_{\mathcal{G}}(\mathbf{x})=\sum_{\forall\omega\in\Omega}\frac{p_{Z}(\mathbf{z})}{\sqrt{det(\mathbf{A_{\omega}}^{T}\mathbf{A_{\omega}})}}$$\mathbb{1}_{z\in\omega}$
For any $\mathbf{z_{1}}\in\omega_{1}$ the sum from the above equation can be ignored, since for all other regions the value would be zero.
Taking negative log and expectation on both sides the conditional entropy becomes
$H(p_{\mathcal{G}}(\mathbf{x_{1}});\mathbf{z}\in\omega_{1})=H(p_{Z}(\mathbf{z_{1}}))+log(\sqrt{det(\mathbf{A_{\omega_{1}}}^{T}\mathbf{A_{\omega_{1}}})})$
For a uniform latent distribution and two regions $\omega_{1}$ and $\omega_{2}$, substituting the second term above with $\psi_{\omega_{1}}$
$H(p_{\mathcal{G}}(\mathbf{x_{1}});\mathbf{z_{1}}\in\omega_{1})$- $H(p_{\mathcal{G}}(\mathbf{x_{2}});\mathbf{z_{2}}\in\omega_{2})=\psi_{\omega_{1}}-\psi_{\omega_{2}}$
Appendix F Broader Impact Statement
Appendix G Extra Figures
Refer to captionIncreasing $\psi$
Figure 17: Local Scaling is sensitive to natural image variations. ImageNet images ordered along the columns (from left to right), with increasing local scaling $\psi$ of the Stable Diffusion decoder learned manifold. We observe that ImageNet samples with lower values of $\psi$ contain simpler backgrounds with modal representation of the object category. Conversely for higher $\psi$ we have increasing diversity both in background and foreground features.
Low local rank, $\nu\downarrow$
High local rank, $\nu\uparrow$
Figure 19: Influence of the local rank descriptor value on overall image perception. Images with the lowest (left) and highest (right) local rank $\nu$ from a set of $20000$ randomly sampled ImageNet dataset samples. Low rank images contain simpler textures for every class compared to the high rank samples. This is because for images with higher local rank, the learned manifold is higher dimensional therefore allowing higher independent degrees of variations locally for the generated images.
Low local scaling, $\psi\downarrow$
High local scaling, $\psi\uparrow$
Figure 20: Influence of the local scaling descriptor.Imagenet images with high and low local scaling for the stable diffusion decoder. Each coordinate in both left and right image grids, correspond to the same imagenet class.